# LORE: Lagrangian-Optimized Robust Embeddings for Visual Encoders

### v1

@@ Abstract
Visual encoders have become fundamental components in modern computer vision pipelines. However, ensuring robustness against adversarial perturbations remains a critical challenge. Recent efforts have explored both supervised and unsupervised adversarial fine-tuning strategies. We identify two key limitations in these approaches: (i) they often suffer from instability, especially during the early stages of fine-tuning, resulting in suboptimal convergence and degraded performance on clean data, and (ii) they exhibit a suboptimal trade-off between robustness and clean data accuracy, hindering the simultaneous optimization of both objectives. To overcome these challenges, we propose Lagrangian-Optimized Robust Embeddings (LORE), a novel unsupervised adversarial fine-tuning framework. LORE utilizes constrained optimization, which offers a principled approach to balancing competing goals, such as improving robustness while preserving nominal performance. By enforcing embedding-space proximity constraints, LORE effectively maintains clean data performance throughout adversarial fine-tuning. Extensive experiments show that LORE significantly improves zero-shot adversarial robustness with minimal degradation in clean data accuracy. Furthermore, we demonstrate the effectiveness of the adversarially fine-tuned CLIP image encoder in out-of-distribution generalization and enhancing the interpretability of image embeddings. The code is available on GitHub.
@@ 1 Introduction
In recent years, embeddings from foundation Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), have transformed downstream tasks, including classification (Conde and Turgutlu, 2021), object detection (Zhong et al., 2021), segmentation (Xu et al., 2022), and image generation (Saharia et al., 2022). These models, trained on large-scale web data, semantically align inputs from different modalities into a joint embedding space. enabling remarkable zero-shot generalization capabilities like zero-shot image classification, where virtually any class can be encoded via its textual description.(Radford et al., 2021; Qian and Hu, 2024). Furthermore, CLIP-like models serve as a critical component in recent large vision language models, such as OpenFlamingo (Awadalla et al., 2023) and LLaVA (Liu et al., 2023).
@@
However, visual encoders, such as CLIP models, exhibit significant vulnerabilities. Adversarial attacks (Goodfellow et al., 2015) and backdoor vulnerabilities (Bai et al., 2024; Liang et al., 2024) can lead to erroneous and misleading embeddings. These challenges highlight the need for robust alignment methods in large, pre-trained visual encoders. To address this, adversarial fine-tuning plays a crucial role in enhancing the robustness of deep neural networks against adversarial examples (Pang et al., 2020; Kurakin et al., 2017; Madry et al., 2019). Moreover, maintaining accuracy and correct embeddings is vital for downstream applications, as it ensures the model’s versatility and reliability without compromising clean data performance, i.e., its nominal performance (Tsipras et al., 2019; Dobriban et al., 2022). While recent works have proposed effective methods to enhance the robustness of visual encoders (Schlarmann et al., 2024; Mao et al., 2023), they often struggle to manage the trade-off between robustness and nominal performance effectively. Therefore, principled approaches are required to explicitly balance this trade-off (Zhang et al., 2019).
@@
In this work, we focus on the vision encoder and introduce Lagrangian-Optimized Robust Embeddings (LORE), a novel unsupervised adversarial fine-tuning framework that effectively balances robustness and nominal performance through a novel constrained optimization framework. While we primarily target CLIP’s image encoder, due to its widespread adoption and strong zero-shot capabilities, we also demonstrate the applicability of LORE to other foundation models such as DINOv2 (Oquab et al., 2024). LORE enforces proximity to a reference model in the embedding space by leveraging the Lagrangian dual method (Bertsekas, 1997), enabling robustness improvements without compromising semantic fidelity on clean data.
@@
Our main contributions are as follows: 1) We propose LORE, a novel unsupervised adversarial fine-tuning framework that optimally balances the trade-off between robustness and nominal performance, with a hyperparameter ($\rho$) enabling principled control over this trade-off. 2) We demonstrate the effectiveness of LORE in image classification, significantly improving zero-shot adversarial robustness over CLIP while maintaining minimal degradation in clean accuracy. 3) We analyze the behavior of the adversarially fine-tuned LORE-CLIP encoder in terms of out-of-distribution generalization and show that it improves the interpretability of image embeddings, highlighting its advantages over other unsupervised adversarial fine-tuning baselines.
@@ 2 Related Work
Adversarial Robustness: Attacks and Defenses. Adversarial attacks exploit neural network vulnerabilities by crafting imperceptible perturbations, leading to misclassifications (Moosavi-Dezfooli et al., 2016; Abdolahpourrostam et al., 2024). Defenses like adversarial training improve robustness but struggle with generalization (Madry et al., 2019). Alternatives such as randomized smoothing and gradient masking face computational overhead or adaptive attack vulnerabilities (Cohen et al., 2019; Athalye et al., 2018). Balancing robustness and nominal performance remains challenging (Tsipras et al., 2019), necessitating more versatile defenses, especially in multimodal settings.
Robustness in Multimodal and VLM Settings. Recent studies highlight the vulnerability of VLMs to adversarial attacks. (Zhang et al., 2024) shows that multimodal embeddings can be disrupted by Adversarial Illusions. TeCOA (Mao et al., 2023) introduces a supervised adversarial training framework for CLIP, improving robustness by aligning adversarial image and text embeddings, but it assumes normalized embeddings and uses simplistic text descriptions, limiting its effectiveness. PMG-AFT (Wang et al., 2024) extends this approach by minimizing the distance between features of adversarial examples and those of the pre-trained model. (Zhou et al., 2024) extend defenses with MMCoA, aligning adversarial and clean embeddings across modalities to protect against various attacks. (Yu et al., 2024) proposed a supervised method that improves zero-shot adversarial robustness of CLIP by aligning text-guided attention maps from clean and adversarial inputs. Among these methods, which rely on labeled data, FARE (Schlarmann et al., 2024) stands out as an unsupervised adversarial fine-tuning strategy that strengthens the CLIP vision encoder. However, FARE lacks explicit mechanisms to balance robustness and task performance. Due to its unsupervised nature, we consider FARE a baseline for evaluating our proposed approach.
Figure 1: (a) clean data accuracy during adversarial fine-tuning with different training perturbation strengths $\varepsilon$, using the loss from Eq. (2). Larger $\varepsilon$ values result in substantial drops in clean data accuracy and early training instability. LORE (with $\rho=0.1$) mitigates this effect, even on ($\varepsilon=10$), maintaining stable and higher clean data accuracy. (b) Pareto frontier comparison between naive regularization of Eq. (2) (blue, varying $\lambda$) and LORE (orange, varying $\rho$). LORE yields a strictly better Pareto front, achieving optimal trade-offs between robustness and clean data accuracy.
@@ 3 Background
Unsupervised Adversarial Fine-Tuning.  FARE (Schlarmann et al., 2024) proposed an unsupervised method for fine-tuning CLIP vision encoder $\phi$. FARE’s goal is to minimize the Euclidean distance between clean and perturbed embeddings under worst-case perturbation conditions:
@@ 5 LORE: Lagrangian-Optimized Robust Embeddings
Our findings in Section 4 highlight the need for a principled method that both stabilizes adversarial fine-tuning and offers a favorable robustness-accuracy trade-off. Constrained learning offers a structured approach to balancing competing objectives like robustness and nominal accuracy. The constrained optimization problem for adversarial robustness in classification, as explored in Robey et al. (2021), can be formulated as:
@@
$\mathrm{s.t.}$ $\quad\ell(f_{\theta}(x),y)\leq\rho,\quad\text{for almost every }(x,y)\in\mathcal{D},$
@@
Extending the optimization framework of (4) to an unsupervised setting, we propose enforcing proximity to a reference model $\phi_{\theta_{0}}$ (e.g., a pre-trained model) in the embedding space $\mathcal{Y}$, thereby eliminating the dependence on $y$. Our formulation is:
@@
$\mathrm{s.t.}$ $\quad d(\phi_{\theta}(x),\phi_{\theta_{0}}(x))\leq\rho\,m(x),\quad\text{for almost every }x\in\mathcal{D}.$
Here, $d:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a divergence metric, and $m:\mathcal{X}\to\mathbb{R}^{+}$ is a sample-specific tolerance margin, independent of $\theta$, used to modulate the constraint across inputs, defining a per-sample margin within which deviations are considered acceptable. For example, when using Euclidean distance, setting $m(x)=\|\phi_{\theta_{0}}(x)\|_{2}$ ensures the constraint becomes scale-invariant with respect to the original embedding magnitude.
Solving the Constrained Problem.  Constrained optimization in deep learning presents significant challenges due to the high-dimensional and non-convex nature of neural networks. Therefore, following (Robey et al., 2021; Chamon et al., 2020; Chamon and Ribeiro, 2021), we leverage Lagrangian duality to obtain approximate solutions for (5) that generalize effectively with respect to both adversarial and nominal performance. Specifically, we solve the dual optimization problem:
@@
where $\lambda_{\omega}:\mathcal{X}\to\mathbb{R}^{+}$ is the dual network parameterized by $\omega$, with outputs made non-negative via Softplus activation, which ensures valid Lagrange multipliers and avoids the sparse gradients of ReLU. This network simplifies the original optimization over all functions $\lambda:\mathcal{X}\to\mathbb{R}^{+}$. By jointly optimizing $\theta$ and $\omega$, our method achieves two key objectives: (i) enhancing the robustness of the model’s embeddings against adversarial examples through the worst-case perturbation $\delta$; and (ii) maintaining proximity to the reference model’s embeddings on clean data, i.e., $\phi_{\theta_{0}}(x)$, to preserve nominal performance. The theoretical gaps and limitations introduced by our formulation and parameterization are detailed in Appendix B.
@@ 6.1 Controlling the Robustness-Accuracy Trade-off
The parameter $\rho$ directly governs the allowed clean embedding deviation, offering a more precise control knob for the robustness-accuracy trade-off than heuristic weightings. While such trade-offs have been extensively studied in supervised adversarial robustness (Zhang et al., 2019; Xiao et al., 2024; Javanmard et al., 2020), they remain underexplored in unsupervised adversarial fine-tuning.
@@
[…]
@@
We empirically evaluate LORE’s control over the robustness–accuracy frontier by training ten models with varying $\rho$. Figure 1(b) shows the resulting Pareto frontier compared to additive-regularization baselines (Eq. 2), and Figure 3 reports clean-data accuracy and adversarial success rate on the test set. These experiments confirm both the practical tightness of our theoretical bound and the superior trade-off achieved by our constrained-optimization framework.
@@
6.2 Out-of-Distribution Robustness and Embedding Interpretability
To assess how unsupervised adversarial fine-tuning affects out-of-distribution (OOD) robustness, we evaluate a ViT-B/32 CLIP model fine-tuned with LORE on ImageNet, using the ImageNet-C (Hendrycks and Dietterich, 2019), which applies common corruptions to simulate distribution shifts. Since adversarial fine-tuning often reduces clean accuracy, it can lead to degraded OOD performance. As shown in Fig. 4(a), FARE2 suffers a substantial drop under corruptions, whereas LORE2 maintains much stronger generalization. In fact, for certain perturbations (e.g., jpeg compression), LORE2 even outperforms the original pretrained CLIP model, underscoring its robustness to distribution shift.
Adversarial training is also known to enhance embedding interpretability and cross-modal alignment (Croce et al., 2025). To measure this effect in the CLIP vision–language space, we generate 150 natural-language templates using GPT-4 (e.g., “This is a photo of {}”) to capture class semantics, then compute the average cosine similarity between each clean image embedding and its corresponding text embedding. Figure 4(b) shows that both FARE2 and LORE2 improve over the pretrained baseline, but LORE2 consistently achieves higher alignment across all training and evaluation $\varepsilon$ settings, demonstrating superior semantic representation.

### v2

@@ Abstract
Visual encoders have become fundamental components in modern computer vision pipelines. However, ensuring robustness against adversarial perturbations remains a critical challenge. Recent efforts have explored both supervised and unsupervised adversarial fine-tuning strategies. We identify two key limitations in these approaches: (i) they often suffer from instability, especially during the early stages of fine-tuning, resulting in suboptimal convergence and degraded performance on clean data, and (ii) they exhibit a suboptimal trade-off between robustness and clean data accuracy, hindering the simultaneous optimization of both objectives. To overcome these challenges, we propose Lagrangian-Optimized Robust Embeddings (LORE), a novel unsupervised adversarial fine-tuning framework. LORE utilizes constrained optimization, which offers a principled approach to balancing competing goals, such as improving robustness while preserving nominal performance. By enforcing embedding-space proximity constraints, LORE effectively maintains clean data performance throughout adversarial fine-tuning. Extensive experiments show that LORE stabilizes training and significantly improves zero-shot adversarial robustness with minimal degradation in clean data accuracy. Furthermore, we demonstrate the effectiveness of the adversarially fine-tuned image encoder in out-of-distribution generalization and enhancing the interpretability of image embeddings. The code is available on GitHub.
@@ 1 Introduction
In recent years, embeddings from foundation Vision-Language Models (VLMs), such as CLIP (radford_learning_2021), have transformed downstream tasks, including classification (conde_clip-art_2021), object detection (zhong_regionclip_2021), segmentation (xu2022simplebaselineopenvocabularysemantic), and image generation (saharia2022photorealistictexttoimagediffusionmodels). These models, trained on large-scale web data, semantically align inputs from different modalities into a joint embedding space. enabling remarkable zero-shot generalization capabilities like zero-shot image classification, where virtually any class can be encoded via its textual description.(radford_learning_2021; qian2024onlinezeroshotclassificationclip).
@@
However, visual encoders, such as CLIP models, exhibit significant vulnerabilities. Adversarial attacks (goodfellow2015explainingharnessingadversarialexamples) and backdoor vulnerabilities (bai_badclip_2024; liang2024revisitingbackdoorattackslarge) can lead to erroneous and misleading embeddings. These challenges highlight the need for robust alignment methods in large, pre-trained visual encoders. To address this, adversarial fine-tuning plays a crucial role in enhancing the robustness of deep neural networks against adversarial examples (pang2020rethinkingsoftmaxcrossentropyloss; kurakin2017adversarialmachinelearningscale; madry2019deeplearningmodelsresistant). Moreover, maintaining accuracy and correct embeddings is vital for downstream applications, as it ensures the model’s versatility and reliability without compromising clean data performance, i.e., its nominal performance (tsipras2019robustnessoddsaccuracy; dobriban2022provabletradeoffsadversariallyrobust). While recent works have proposed effective methods to enhance the robustness of visual encoders (schlarmann2024robustclip; mao2023understandingzeroshotadversarialrobustness), they often struggle to manage the trade-off between robustness and nominal performance effectively. Therefore, principled approaches are required to explicitly balance this trade-off (pmlr-v97-zhang19p).
@@
In this work, we focus on the vision encoder and introduce Lagrangian-Optimized Robust Embeddings (LORE), a novel unsupervised adversarial fine-tuning framework that effectively balances robustness and nominal performance through a novel constrained optimization framework. While we primarily target CLIP’s image encoder, due to its widespread adoption and strong zero-shot capabilities, we also demonstrate the applicability of LORE to other foundation models such as DINOv2 (oquab2024dinov2learningrobustvisual). LORE enforces proximity to a reference model in the embedding space by leveraging the Lagrangian dual method (Bertsekas1997), enabling robustness improvements without compromising semantic fidelity on clean data.
@@
Our main contributions are as follows: 1) We propose LORE, a novel unsupervised adversarial fine-tuning framework that achieves a superior empirical balance on the trade-off between robustness and nominal performance, with a hyperparameter ($\rho$) enabling principled control over this trade-off. 2) We demonstrate the effectiveness of LORE in image classification, significantly improving zero-shot adversarial robustness over CLIP while maintaining minimal degradation in clean accuracy. 3) We analyze the behavior of the adversarially fine-tuned LORE-CLIP encoder in terms of out-of-distribution generalization and show that it improves the interpretability of image embeddings, highlighting its advantages over other unsupervised adversarial fine-tuning baselines.
@@ 2 Related Work
Adversarial Robustness: Attacks and Defenses. Adversarial attacks exploit neural network vulnerabilities by crafting imperceptible perturbations, leading to misclassifications (goodfellow2015explainingharnessingadversarialexamples; szegedy2014intriguingpropertiesneuralnetworks). Defenses like adversarial training improve robustness but struggle with generalization (madry2019deeplearningmodelsresistant). Balancing robustness and nominal performance remains challenging (tsipras2019robustnessoddsaccuracy; zhang2019theoreticallyprincipledtradeoffrobustness), necessitating more versatile defenses, especially in multimodal settings. A principled approach to managing the trade-off is Constrained Optimization: (robey2021adversarialrobustnesssemiinfiniteconstrained) pioneered using a Lagrangian duality framework to enforce explicit constraints, avoiding heuristic loss weighting.
Robustness in Foundation Encoders. Recent studies have highlighted the vulnerability of foundation visual encoders, e.g. CLIP and DINOv2, to adversarial attacks, demonstrating that embeddings can be disrupted (zhang2024adversarialillusionsmultimodalembeddings). To address this, supervised defenses like TeCoA (mao2023understandingzeroshotadversarialrobustness) and MMCoA (zhou2024revisitingadversarialrobustnessvision) improve robustness by aligning adversarial and clean embeddings, often relying on auxiliary information like text, using simplistic text descriptions, or attention maps,(NEURIPS2024_aec2dfc4), assuming normalized embeddings. PMG-AFT (10656801) extends this approach by minimizing the distance between adversarial features and those of the pre-trained model. In the label-free setting, Unsupervised Adversarial Fine-Tuning seeks to achieve robustness by modifying contrastive losses, leading to methods like Robust Contrastive Learning (RoCL) (kim2020adversarialselfsupervisedcontrastivelearning) and the decoupled approach of DeACL (zhang2022decoupledadversarialcontrastivelearning). Among unsupervised fine-tuning strategies for the CLIP vision encoder, FARE (schlarmann2024robustclip) stands out as a direct baseline. However, FARE and similar methods often lack explicit mechanisms to balance robustness and task performance, resulting in a suboptimal trade-off. Due to its unsupervised nature, we consider FARE a baseline for evaluating our proposed approach.
Figure 1: (a) clean data accuracy during adversarial fine-tuning with different training perturbation strengths $\varepsilon$, using the loss from Eq. (2). Larger $\varepsilon$ values result in substantial drops in clean data accuracy and early training instability. LORE (with $\rho=0.1$) mitigates this effect, even on ($\varepsilon=10$), maintaining stable and higher clean data accuracy. (b) Pareto frontier comparison between naive regularization of Eq. (2) (blue, varying $\lambda$) and LORE (orange, varying $\rho$). LORE yields a strictly better empirical Pareto front, demonstrating superior trade-offs.
@@ 3 Background
Unsupervised Adversarial Fine-Tuning.  FARE (schlarmann2024robustclip) proposed an unsupervised method for fine-tuning CLIP vision encoder $\phi$. FARE’s goal is to minimize the Euclidean distance between clean and perturbed embeddings under worst-case perturbation conditions:
@@ 5 LORE: Lagrangian-Optimized Robust Embeddings
Our findings in Section 4 highlight the need for a principled method that both stabilizes adversarial fine-tuning and offers a favorable robustness-accuracy trade-off. Constrained learning offers a structured approach to balancing competing objectives like robustness and nominal accuracy. The constrained optimization problem for adversarial robustness in classification, as explored in robey2021adversarialrobustnesssemiinfiniteconstrained, can be formulated as:
@@
$\mathrm{s.t.}$ $\qquad\ell(f_{\theta}(x),y)\leq\rho,\qquad\text{for almost every }(x,y)\in\mathcal{D},$
@@
Extending the optimization framework of (4) to an unsupervised setting, we propose enforcing proximity to a reference model $\phi_{\theta_{0}}$ (e.g., a pre-trained model) in the embedding space $\mathbb{R}^{k}$, where $k$ denotes the embedding dimension, thereby eliminating the dependence on $y$. Our formulation is:
@@
$\mathrm{s.t.}$ $\qquad d(\phi_{\theta}(x),\phi_{\theta_{0}}(x))\leq\rho\,m(x),\qquad\text{for almost every }x\in\mathcal{D}.$
Here, $d:\mathbb{R}^{k}\times\mathbb{R}^{k}\to\mathbb{R}^{+}$ is a divergence metric, and $m:\mathcal{X}\to\mathbb{R}^{+}$ is a sample-specific tolerance margin, independent of $\theta$, used to modulate the constraint across inputs, defining a per-sample margin within which deviations are considered acceptable. For example, when using Euclidean distance, setting $m(x)=\|\phi_{\theta_{0}}(x)\|_{2}$ ensures the constraint becomes scale-invariant with respect to the original embedding magnitude.
Solving the Constrained Problem.  Constrained optimization in deep learning presents significant challenges due to the high-dimensional and non-convex nature of neural networks. Therefore, following (robey2021adversarialrobustnesssemiinfiniteconstrained; Chamon2020TheED; chamon2021probablyapproximatelycorrectconstrained), we leverage Lagrangian duality to obtain approximate solutions for (5) that generalize effectively with respect to both adversarial and nominal performance. Specifically, we solve the dual optimization problem:
@@
where $\lambda_{\omega}:\mathcal{X}\to\mathbb{R}^{+}$ is the dual network parameterized by $\omega$, with outputs made non-negative via Softplus activation, which ensures valid Lagrange multipliers and avoids the sparse gradients of ReLU. This network simplifies the original optimization over all functions $\lambda:\mathcal{X}\to\mathbb{R}^{+}$. By jointly optimizing $\theta$ and $\omega$, our method achieves two key objectives: (i) enhancing the robustness of the model’s embeddings against adversarial examples through the worst-case perturbation $\delta$; and (ii) maintaining proximity to the reference model’s embeddings on clean data, i.e., $\phi_{\theta_{0}}(x)$, to preserve nominal performance. Note that this adaptive framework fundamentally differs from regularization, and is automatically enforced if embeddings diverge, The theoretical gaps and limitations introduced by our formulation and parameterization are detailed in Appendix B.
@@ 6.1 Controlling the Robustness-Accuracy Trade-off
The parameter $\rho$ directly governs the allowed clean embedding deviation, offering a more precise control knob for the robustness-accuracy trade-off than heuristic weightings. While such trade-offs have been extensively studied in supervised adversarial robustness (pmlr-v97-zhang19p; xiao2024bridginggaprademachercomplexity; javanmard2020precisetradeoffsadversarialtraining), they remain underexplored in unsupervised adversarial fine-tuning.
@@
Assume $d(\cdot,\cdot)$ is a nonnegative metric on the embedding space. Moreover, for any fixed reference embedding $c$, the map $u\mapsto d(u,c)$ is $L_{d}$-Lipschitz in its first argument, i.e., $\bigl|d(u_{1},c)-d(u_{2},c)\bigr|\leq L_{d}\,\|u_{1}-u_{2}\|$ for some norm $\|\cdot\|$.
@@
We empirically evaluate LORE’s control over the robustness–accuracy frontier by training ten models with varying $\rho$. Figure 1(b) shows the resulting Pareto frontier compared to additive‐regularization baselines (Eq. 2), and Figure 3 reports clean-data accuracy and adversarial success rate on the test set. These experiments confirm both the practical tightness of our theoretical bound and the superior trade‐off achieved by our constrained‐optimization framework.
@@
6.2 Out‐of‐Distribution Robustness and Embedding Interpretability
To assess how unsupervised adversarial fine-tuning affects out-of-distribution (OOD) robustness, we evaluate a ViT-B/32 CLIP model fine-tuned with LORE on ImageNet, using the ImageNet-C (hendrycks2019benchmarkingneuralnetworkrobustness), which applies common corruptions to simulate distribution shifts. Since adversarial fine-tuning often reduces clean accuracy, it can lead to degraded OOD performance. As shown in Fig. 4(a), FARE2 suffers a substantial drop under corruptions, whereas LORE2 maintains much stronger generalization. In fact, for certain perturbations (e.g., jpeg compression), LORE2 even outperforms the original pretrained CLIP model, underscoring its robustness to distribution shift.
Adversarial training is also known to enhance embedding interpretability and cross-modal alignment (croce2024adversarially). To measure this effect in the CLIP vision–language space, we generate 150 natural‐language templates using GPT-4 (e.g., “This is a photo of {}”) to capture class semantics, then compute the average cosine similarity between each clean image embedding and its corresponding text embedding. Figure 4(b) shows that both FARE2 and LORE2 improve over the pretrained baseline, but LORE2 consistently achieves higher alignment across all training and evaluation $\varepsilon$ settings, demonstrating superior semantic representation.